The LD 50 for a substance in the dose that would be lethal for 50% of the population that receive that does the LD50 for aspirin in rats is 1.75 g per kilogram of body weight . how many aspirin tablets each of which contains 325 mg of aspirin would you have to feed a rat weigh 290 g to give him that dose?


1
Expert's answer
2019-09-09T04:57:54-0400

If 1.75 g of aspirin is LD50 for a kilo of rat body weight, then find LD50 for 290 g rat:

"m = 1.75*0.29 = 0.5075 (g)"

0.5075 g or 507.5 mg of aspirin is LD50 for a rat weigh 290 g.

Find how many tablets contain 507.5 mg of aspirin:

"n = 507.5\/325 = 1.56 (u.)"

Answer: 1.56 tablets of aspirin contain LD50 for a rat weigh 290 g.